# TreeLearn: A Comprehensive Deep Learning Method for Segmenting Individual Trees from Ground-Based LiDAR Forest Point Clouds

### v2

@@ 4.2 Clean Labels
Following the same procedure as described in Section 4.1, $4500$ training crops were generated for the cleanly labeled forest plot. For fine-tuning, the model was initialized with parameters obtained by training with noisy labels. From there, the network was trained for 3125 iterations without warm up and a maximum/minimum learning rate of $4\times 10^{-5}$/$5\times 10^{-7}$. Otherwise, the training specifications were the same as described in Section 4. Fine-tuning was performed on a single Nvidia RTX A5000 and took roughly 2 hours. Apart from fine-tuning, training the model from scratch with cleanly labeled data was also attempted, but was not successful due to overfitting.
4.3 Model Selection
We had one cleanly labeled forest plot. To report unbiased estimates of model performance, we split it into two parts: one used for model selection (validation set) and the other one for estimating model performance (test set). To make the best use of the available data, we used each part once as the validation set and evaluated the performance on the other, and vice versa. When training with the noisy labels, the parameters with the lowest offset loss on the validation set were selected. The validation set was then also used to fine-tune the selected parameters as described in Section 4.2. In this case, the parameters after the last training iteration were selected without any selection criterion. This way, test performance on the full cleanly labeled forest plot can be evaluated for both noisy and clean label training.
@@ 4.3 Model Selection
For the density based clustering step described in Section 3.4, there are four hyperparameters. Two hyperparameters decide which points are considered in the clustering algorithm: the maximum vertical distance to the tree base was set to $\tau_{\textrm{off}}=2~{}\text{m}$ and the minimum required verticality feature value was set to $\tau_{\textrm{vert}}=0.6$. The verticality feature of a point can reach a maximum value of one for a perfectly vertical straight line of points. A value of $0.6$ allows for slightly slanted trunks. A value of 2 m for the maximum vertical distance from the tree base further ensures that only trunk points are considered. There are two additional hyperparameters that govern the behavior of the clustering algorithm: first, the minimum number of points for a valid cluster was set to $\tau_{\textrm{min}}=100$. This number was chosen based on considerations how many points a valid tree cluster might contain in extreme cases, e.g. when the tree trunk is very thin. Second, the grouping radius for clustering was set to $\tau_{\textrm{group}}=0.15~{}\text{m}$. In Appendix A, we analyze how the number of false positive and false negative predictions is influenced by different grouping radii $\tau_{\textrm{group}}$. Together with qualitative results obtained on an out-of-domain test dataset (see Section 6.2), this analysis suggests that our method can be employed out-of-the-box without extensive hyperparameter tuning.
5 Evaluation Metrics
5.1 Semantic Segmentation Evaluation: Tree vs. Non-Tree
For semantic segmentation evaluation, we assess how well the method divides the points into the semantic classes tree and non-tree. As a metric, we choose the accuracy, which is the number of correctly predicted points divided by the total number of points. We report evaluation results on the point cloud subsampled with a voxel size of $10\text{cm}^{3}$ (see Section 3.1). This avoids a disproportionate weighting of regions with higher point densities like the trunks.
5.2 Instance Detection Evaluation
For instance detection evaluation, we are interested in how well the predicted tree instances match the ground truth tree instances. Specifically, we consider true positive predictions ($TP$), false negative predictions ($FN$) and false positive predictions ($FP$) as defined by [14]: a tree prediction is $TP$ if, in principle, it has a clear match (defined below) to a ground truth tree. If a tree is predicted to belong to a larger nearby tree or is not detected at all, it is $FN$. Finally, if a prediction is no complete tree instance but only a fragment, it is $FP$. It should be noted that one prediction consisting of two merged trees only results in one $FN$. The larger tree of the two still counts as detected using the above definition.
To automatically identify $TP$, $FP$ and $FN$ predictions, we first calculate the pointwise intersection over union (IoU) between all predictions and ground truths. Then, Hungarian matching [63] is performed to assign predictions to ground truths. To avoid the assignment of fragments to ground truth trees, we set a minimum IoU of 0.3 as a requirement for a match. Matched predictions, non-matched predictions and non-matched ground truths are considered $TP$, $FP$ and $FN$ respectively. Trees at the edge of the segmented area are often only partially within the point cloud and might consist of only one branch within the segmented area. If this one branch is falsely assigned to another tree, this results in a $FN$. This is misleading since the actual tree, which is not within the evaluation area, might be correctly detected. Therefore, we consider only trees that are fully within the prediction area (156 trees) for this part of the evaluation. Likewise, only predictions whose maximum overlap is with one of these 156 trees, are taken into account.
5.3 Instance Segmentation Evaluation
@@
For the evaluation of instance segmentation, we assess how well the points of the predicted tree instances match the points of the corresponding ground truth tree instances. This is different to instance detection (Section 5.2), in that we are not evaluating whether a specific tree is detected at all, but how well the individual points of a specific predicted tree match to the actual tree. Since detection errors represent a different quality of mistakes, we do not consider predicted tree instances that are associated with a detection error. This includes (1) $FP$ predicted tree instances (fragments) and (2) predicted tree instances that consist of more than one ground truth tree (merged predictions). To evaluate the instance segmentation, we use precision, recall and the F1-score:
$Prec=\frac{TP}{TP+FP}\quad\qquad Rec=\frac{TP}{TP+FN}\quad\qquad F1=2*\frac{Prec*Rec}{Prec+Rec}$ (1)
Note that $TP$, $FP$ and $FN$ are here defined at the point-level, not the instance-level. The metrics are calculated separately for every predicted tree instance with the corresponding ground truth tree points as positives and all other points as negatives. Then, they are averaged over all predictions that are taken into account.
@@
Not all locations of a tree are equally difficult to segment correctly. For example, points near the trunk are usually easier to assign than points farther outside where there are many interactions with other trees. To quantify how well different parts of the trees are segmented, we partition the points of a prediction and the corresponding tree into 10 subsets and calculate the metric for each subset. For each subset, the metric is then averaged across predictions. We propose two axes for partitioning: (1) horizontal distance to the trunk, and (2) vertical distance to the forest ground. For the horizontal partition, the $i$-th subset contains all points with a horizontal distance to the trunk between $\frac{i-1}{10}r$ and $\frac{i}{10}r$ where $r$ is the maximum distance. For the vertical partition, the $i$-th subset contains all points with a vertical distance to the ground between $\frac{i-1}{10}h$ and $\frac{i}{10}h$ where $h$ is the height of the tree. We set $i=1,\ldots,10$.
6 Results
Figure 4: Comparison of instance segmentation results for two heavily intertwined trees. Areas of interest are marked with red circles. The result of TreeLearn has some errors, but is better than Lidar360.
Table 1: Overall results. Instance segmentation results averaged across predictions. Semantic and instance segmentation results in %.
Semantic Segmentation Detection Instance Segmentation
Accuracy $FP$ Predictions $FN$ Trees $F1$
TLS2treesa 98.9 0 0 88.39
Lidar360 99.6 1 1 93.23
Noisy Labels Training 99.69 0 0 93.98
Clean Labels Fine-tuning 99.88 1 0 97.84
newaTLS2trees detects very small tree-like formations. For a fair comparison, these detections were manually assigned to the
newt non-tree class since we only consider trees of at least 10 m height.
TreeLearn performs three tasks: separating tree points from non-tree points, detecting individual trees and assigning points to their corresponding tree instance. In the following, we evaluate the system’s performance on each of these tasks using our manually labeled benchmark dataset and compare it with two existing methods: the commercial Lidar360 software [50] and TLS2trees [33], a recent open source graph-based instance segmentation method.
We trained TreeLearn supervised on labels generated automatically by Lidar360. The segmentation results of this software have mistakes, hence the need to improve. Our first, perhaps surprising, finding is that the segmentation algorithm learned from these ‘noisy’ labels performs slightly better than the algorithm that generated the training data, i. e. a machine learning approach can extract more consistent segmentation rules than its ‘teacher’. On average, TreeLearn improves instance segmentation performance measured by the F1-score from 93.23 % (Lidar360) to 93.98 % (Table 1). TLS2trees achieves an F1-score of 88.39 %. TLS2trees was applied using the standard hyperparameters from the paper [33] to give an example of the out-of-the-box performance of another method. It should be noted that TLS2trees, unlike our model, has not been developed using the specific forest type and laser scanner used in this study. We expect the reported performance gap between the two methods to decrease for other settings.
Fine-tuning TreeLearn on the clean manually labeled forest plot improves the result drastically to an instance segmentation performance of 97.84 % (for an example, see Figure 4). Both variations of TreeLearn also produce a better separation of trees and non-tree points than Lidar360 and TLS2trees. In particular, the semantic segmentation accuracy increases from 99.6 % (Lidar360) to 99.88 % for the fine-tuned TreeLearn. Detection performance on our benchmark dataset is without errors for both TLS2trees and TreeLearn trained with noisy labels. In contrast, the fine-tuned TreeLearn and Lidar360 produce one prediction that does not actually constitute a tree ($FP$) and Lidar360 additionally fails to detect one tree ($FN$). However, detection errors are too rare on our benchmark dataset, so there is no conclusive evidence that one method performs better than the others in this regard.
@@ 6.1 Instance Segmentation
Table 2: Instance segmentation results averaged across predictions. All values in %.
Precision Recall $F1$
TLS2trees 88.2 90.52 88.39
Lidar360 92.56 94.82 93.23
Noisy Labels Training 93.45 95.34 93.98
Clean Labels Fine-tuning 97.57 98.28 97.84
In this section, we take a more detailed look at the instance segmentation results. In addition to an overall better F1-score, both TreeLearn settings achieve a better precision and recall value than Lidar360 and TLS2trees (Table 2). We also observe that, irrespective of the method, recall is higher than precision. We conjecture that this is the case because the branches of larger trees often protrude into smaller trees. If these branches are then incorrectly assigned to the smaller tree, the precision of the smaller tree is affected considerably. On the other hand, a falsely assigned branch does not affect the recall of the larger tree to the same degree.
@@
Figure 5: Instance segmentation performance. (a) shows metrics for segments along the x- and y- axis. (b) shows metrics for segments along the z-axis.
Figure 6: Comparison of semantic segmentation results. Green points are assigned to the non-tree class, red points to the tree class. Lidar360 assigns completely unrelated points to the tree class. TreeLearn is more accurate.
Figure 7: Segmentation result generated by TreeLearn on a point cloud from Wytham Woods forest.
@@
A more fine-grained understanding can be achieved by evaluating the segmentation performance at different locations of a tree. For this, we partition the points of a prediction and the corresponding tree into subsets based on horizontal distance to the tree trunk and vertical distance to the forest ground as described in Section 5.3. Figure 5 shows precision, recall and F1-scores for the two partitions. For all methods, performance decreases for points farther away from the trunk due to increased interactions with other trees. However, this trend is less pronounced for TreeLearn, indicating a certain stability with respect to the distance to the trunk. This is especially true for the fine-tuned model.
For the vertical partition of the trees, both TreeLearn and Lidar360 have a visible performance low for intermediate heights, where trees tend to have their maximum width. For TLS2Trees, this tendency is not observed. Furthermore, the area directly above the ground poses difficulties for all methods. This is most likely due to understory and tree points that are hard to distinguish from each other (see e.g. Figure 6).
6.2 Qualitative Test Results
In addition to the quantitative results on our benchmark dataset, we used our fine-tuned TreeLearn model to obtain qualitative results on a completely unrelated forest plot. Specifically, we used a forest point cloud from the Wytham Woods forest in Oxford, UK, that was kindly provided by the authors that introduced the dataset [64, 10]. This forest point cloud differs in the tree species composition and was scanned with a different laser scanner than our training data. Our method produces adequate segmentation results on this out-of-domain forest point cloud without any further hyperparameter tuning (Figure 7).

### v3

@@ 2.3.2 Fine-tuning
Following the same procedure as described in Section 2.3.1, 30 0003000030\,00030 000 training crops in total were generated for the three datasets. To quantify the effect of using more training data, three fine-tuning settings were employed that differ in the datasets used (Table 2). The small setting uses the same 42 training plots as the conceptually similar ForAINet [41], allowing a direct comparison. The mid setting was additionally trained on the validation and test data of For-Instance and the LAUTx dataset. For the large setting, we additionally trained on the Wytham Woods plot. In all settings, the model was initialized with parameters obtained by training with noisy labels. From there, the network was trained for 12 5001250012\,50012 500 iterations without warm-up and a maximum/minimum learning rate of $5\times 10^{-4}$/$5\times 10^{-5}$. Otherwise, the training specifications were the same as described in Section 2.3. Fine-tuning was performed on a single Nvidia RTX A5000 and took roughly 3 hours. A single held-out plot from NIBIO (plot 16) was used as the validation set for all settings. The model with the best offset loss on the validation set was selected as the final model for benchmarking the model performance.
Table 2: Fine-tuning settings. Checkmarks indicate that the respective dataset was used for fine-tuning.
Name For-Instance train For-Instance val/test LAUTx Wytham Woods L1W
noisy - - - - -
small - - - -
mid - -
large -
2.4 Hyperparameter selection
@@
For the density based clustering step described in Section 2.2.4, there are four hyperparameters. Two hyperparameters decide which points are considered in the clustering algorithm: the maximum vertical distance to the tree base was set to $\tau_{\textrm{off}}=2~{}\text{m}$ and the minimum required verticality feature value was set to $\tau_{\textrm{vert}}=0.6$. The verticality feature of a point can reach a maximum value of one for a perfectly vertical straight line of points. A value of $0.6$ allows for slightly slanted trunks. A value of 2 m for the maximum vertical distance from the tree base further ensures that primarily trunk points are considered. There are two additional hyperparameters that govern the behavior of the clustering algorithm: first, the minimum number of points for a valid cluster was set to $\tau_{\textrm{min}}=100$. This number was chosen based on considerations how many points a valid tree cluster might contain in extreme cases, e.g. when the tree trunk is very thin. Second, the grouping radius for clustering was set to $\tau_{\textrm{group}}=0.15~{}\text{m}$. In Appendix C, we analyze how the number of commission and omission errors is influenced by different grouping radii $\tau_{\textrm{group}}$. This analysis suggests that our method can be employed out-of-the-box without extensive hyperparameter tuning.
2.5 Evaluation metrics
For evaluating the quality of segmentation results, we are interested in three aspects: Separating tree points from non-tree points (semantic segmentation), detecting individual trees (instance detection) and assigning points to their corresponding tree instance (instance segmentation). As mentioned in Section 2.1.2, small trees are often labeled inconsistently in the existing datasets. Therefore, semantic segmentation results will be greatly influenced by which training data is used and which dataset is used for evaluation. At the same time, all of the methods compared in this work achieve a reasonable segmentation into tree and non-tree points which is why we decided to perform no evaluation of this aspect. For instance detection and instance segmentation evaluation we employ standard metrics as described in the following Sections 2.5.1 and 2.5.2. For all evaluations done in this work, we subsampled the ground truth point clouds with a voxel size of 10 cm. This is recommended since it avoids a disproportionate weighting of regions with higher point densities. For example, the trunks of ground-based scans usually have a much higher density than the tree crowns. Instructions and code for running our evaluation on arbitrary forest point clouds is publicly available in the code base of this work.
2.5.1 Instance detection evaluation
For instance detection evaluation we denote the set of ground truth tree instances as $\left\{I_{\text{i}}^{\text{gt}},i\in\{1,\dots,N_{\text{gt}}\}\right\}$ and the set of predicted tree instances as $\left\{I_{\text{j}}^{\text{pred}},j\in\{1,\dots,N_{\text{pre}}\}\right\}$. For each ground truth tree and each predicted tree, we first calculate the pointwise intersection over union (IoU):
$\text{IoU}(I_{\text{i}}^{\text{gt}},I_{\text{j}}^{\text{pred}})=\frac{TP_{\text{ij}}}{TP_{\text{ij}}+FP_{\text{ij}}+FN_{\text{ij}}}$
where $TP_{\text{ij}}$, $FP_{\text{ij}}$ and $FN_{\text{ij}}$ denote the number of true positive, false positive and false negative points for $I_{\text{i}}^{\text{gt}}$ and $I_{\text{j}}^{\text{pred}}$. We obtain an IoU-matrix of size $N_{\text{gt}}\times N_{\text{pre}}$. With this IoU-matrix we perform the Hungarian algorithm [70] to match ground truths with predictions. Following Xiang et al. [41], we then filter out matches with an IoU smaller than 0.5 to obtain the final number of matched ground truths $N_{\text{gt}}^{\text{m}}$, unmatched ground truths $N_{\text{gt}}^{\text{u}}$ and unmatched predictions $N_{\text{pred}}^{\text{u}}$. These values are then used to define the metrics completeness, omission error, commission error, and F1-score:
$C=\frac{N_{\text{gt}}^{\text{m}}}{N_{\text{gt}}^{\text{m}}+N_{\text{gt}}^{\text{u}}}=\frac{N_{\text{gt}}^{\text{m}}}{N_{\text{gt}}}$
$E_{\text{om}}=\frac{N_{\text{gt}}^{\text{u}}}{N_{\text{gt}}^{\text{m}}+N_{\text{gt}}^{\text{u}}}=1-C$
$E_{\text{com}}=\frac{N_{\text{pred}}^{\text{u}}}{N_{\text{gt}}^{\text{m}}+N_{\text{pred}}^{\text{u}}}=\frac{N_{\text{pred}}^{\text{u}}}{N_{\text{pred}}}$
$F\text{1}=\frac{2\cdot(1-E_{\text{om}})\cdot(1-E_{\text{com}})}{(1-E_{\text{om}})+(1-E_{\text{com}})}$
Many forest datasets have unlabeled trees at the edges. These trees are not removed during evaluation because they potentially contribute to segmentation errors, which should be taken into account. For example, part of an unlabeled tree might be falsely assigned to a prediction for a labeled ground truth tree. Apart from unlabeled trees at the edges, small trees are often not labeled. It would be inadequate to count model predictions as unmatched if they correspond to such unlabeled trees. Therefore, unmatched predictions are removed if less than half of their points belong to a labeled tree. This way, only unmatched predictions that are associated with a labeled ground truth tree are taken into account, which represent critical errors such as oversegmentation.
2.5.2 Instance segmentation evaluation
@@
For the evaluation of instance segmentation, we want to assess how well the points of the predicted tree instances match the points of the corresponding ground truth tree instances. This is different from instance detection (Section 2.5.1), in that we are not evaluating whether a specific tree is detected at all, but how precisely each ground truth tree is segmented. We follow the evaluation protocol of Xiang et al. [41]. For each ground truth tree $I_{\text{i}}^{\text{gt}}$, we determine the index $max_{\text{i}}$ of the predicted tree with the highest IoU score:
$max_{\text{i}}=\operatornamewithlimits{argmax}_{j=1}^{N_{\text{pre}}}\left(\text{IoU}\left(I_{i}^{\text{gt}},I_{j}^{\text{pre}}\right)\right)$
This index is used to pair ground truths with predictions. It should be noted that different ground truths can be associated with the same prediction this way. For example, if a prediction is a merged tree that consists of two ground truth trees, then both of these ground truth trees will be associated with this prediction. This pairing protocol ensures that every ground truth tree can be associated with a prediction. This enables a systematic comparison of segmentation performance even if the number of matched trees according to the criterion defined in Section 2.5.1 differs between methods. Following Xiang et al. [41], we employ the coverage which is defined as the mean IoU across all pairs of ground truth and prediction:
$Cov=\frac{1}{N_{\text{gt}}}\sum_{i=1}^{N_{\text{gt}}}\text{IoU}(I_{\text{i}}^{\text{gt}},I_{max_{\text{i}}}^{\text{pred}})$
For a more detailed performance evaluation, we also calculate the mean precision and recall:
$Prec=\frac{1}{N_{\text{gt}}}\sum_{i=1}^{N_{\text{gt}}}\frac{TP_{\text{i},max_{\text{i}}}}{TP_{\text{i},max_{\text{i}}}+FP_{\text{i},max_{\text{i}}}}$
$Rec=\frac{1}{N_{\text{gt}}}\sum_{i=1}^{N_{\text{gt}}}\frac{TP_{\text{i},max_{\text{i}}}}{TP_{\text{i},max_{\text{i}}}+FN_{\text{i},max_{\text{i}}}}$
@@
Not all locations of a tree are equally difficult to segment correctly. For example, points near the trunk are usually easier to assign than points farther outside where there are many interactions with other trees. To quantify how well different parts of the trees are segmented, we partition the points of a ground truth tree and the corresponding prediction into ten subsets and calculate the metric for each subset. For each subset, the metric is then averaged across predictions. We propose two axes for partitioning: (1) horizontal distance to the trunk, and (2) vertical distance to the forest ground. For the horizontal partition, the $i$-th subset contains all points with a horizontal distance to the ground truth trunk between $\frac{i-1}{10}r$ and $\frac{i}{10}r$ where $r$ is the ground truth maximum distance. For the vertical partition, the $i$-th subset contains all points with a vertical distance to the ground between $\frac{i-1}{10}h$ and $\frac{i}{10}h$ where $h$ is the ground truth height of the tree. We set $i=1,\ldots,10$. It should be noted that points of a prediction that are farther away than $r$ and $h$ from the trunk and ground, respectively, are not taken into account in this part of the evaluation.
3 Results and dicsussion
TreeLearn takes unlabeled forest point clouds as input, identifies tree points and divides them into individual instances. We evaluate how accurately TreeLearn detects and segments trees on our benchmark dataset L1W and the Wytham Woods dataset, and explore whether performance improves with additional training data. We also provide a comparison with ForAINet [41] and SegmentAnyTree [42] that are conceptually similar to TreeLearn. The results of these methods were obtained via a cloud-based service (https://forestsens.com/, 26th Aug 2024). In addition, we report results of the Lidar360 algorithm [53] and TLS2Trees [37] on L1W to provide a comparison with classical algorithms based on hand-crafted features. TLS2Trees was employed with the standard hyperparameters from their work [37].
3.1 Instance detection
Regarding the instance detection results on L1W (Table 3), TreeLearn trained with noisy labels achieves the lowest omission and commission error rates ($E_{\text{om}}$=0.6%, $E_{\text{com}}$=0.6%), closely matching the performance of Lidar360 that was used to generate the noisy labels ($E_{\text{om}}$=1.3%, $E_{\text{com}}$=1.3%). Interestingly, the number of commission errors on L1W increases through fine-tuning on manually labeled data (e.g. $E_{\text{com}}$=4.3for TreeLearn large). In Section 3.4, we provide a detailed error analysis for TreeLearn, where we discuss this finding. ForAINet ($E_{\text{om}}$=15.4%, $E_{\text{com}}$=27.1%) and SegmentAnyTree ($E_{\text{om}}$=10.9%, $E_{\text{com}}$=49.1%) exhibit the highest error rates. We conjecture that this is the result of the block-merging algorithm these methods rely on: Instance predictions are first obtained on cylinders with a radius of 8 m and subsequently merged based on an IoU-threshold. Inadequate merging could lead to over- and undersegmentation of trees (Figure 4b and c). In contrast, TreeLearn generates tree clusters for the whole input plot in a single step, leading to more accurate tree instances (Figure 4) even when trained on the same training data as ForAINet (TreeLearn small).
Table 3: Quantitative results in %.
Instance detection Instance segmentation
Test dataset Method $C$ $E_{\text{om}}$ $E_{\text{com}}$ $F\text{1}$ $Prec$ $Rec$ $Cov$
L1W ForAINet 84.6 15.4 27.1 78.3 84.1 86.1 73.8
SegmentAnyTree 89.1 10.9 49.1 64.8 90.2 85.2 77.5
TLS2Trees 96.2 3.8 4.5 95.8 87.9 90.5 79.9
Lidar360 98.7 1.3 1.3 98.7 91.7 94.5 87.0
TreeLearn noisy 99.4 0.6 0.6 99.4 92.0 94.2 86.7
TreeLearn small 99.4 0.6 4.9 97.2 93.3 94.9 88.8
TreeLearn mid 100.0 0.0 4.3 97.8 94.8 96.2 91.3
TreeLearn large 99.4 0.6 4.3 97.5 95.1 96.3 91.8
Wytham ForAINet 36.8 63.2 42.9 44.8 44.5 84.2 39.7
Woods SegmentAnyTree 39.0 61.0 52.8 42.7 47.6 80.8 41.2
TreeLearn noisy 39.9 60.1 32.2 50.2 41.2 71.3 37.9
TreeLearn small 47.9 52.1 31.0 56.5 49.5 90.9 46.1
TreeLearn mid 52.9 47.1 26.0 61.7 52.5 91.9 49.2
On the more challenging Wytham Woods dataset, performance is generally worse. TreeLearn noisy performs worst here. This is unsurprising as trees smaller than 10 m were labeled as non-tree in the training data. Wytham Woods contains many such small trees that are not detected by the model. Fine-tuning on manually labeled data that contains trees smaller than 10 m resolves this property. Nonetheless, the performance gap between TreeLearn small ($E_{\text{om}}$=52.1%, $E_{\text{com}}$=31.0%) and ForAINet ($E_{\text{om}}$=63.2%, $E_{\text{com}}$=42.9%) and SegmentAnyTree ($E_{\text{om}}$=61.0%, $E_{\text{com}}$=52.8%) is smaller than on L1W. This is most likely due to the presence of many extremely difficult trees, for example with very closely spaced trunks (Figure 4a), where both models perform equally poorly.
@@ 3.2 Instance segmentation
TreeLearn achieves a superior instance segmentation performance compared to the other methods (Table 3). When training TreeLearn only with labels automatically generated by Lidar360 (TreeLearn noisy), our method closely matches the metrics of Lidar360. This demonstrates that the network successfully learned the segmentation rules of the algorithm. When fine-tuning TreeLearn with the same data used to train ForAINet (TreeLearn small), the coverage increases for both L1W (88.8%) and Wytham Woods (46.1%). On L1W, this is the case even though the number of commission errors increases, which negatively impacts the coverage. ForAINet only achieves a coverage of 73.8% and 39.7% on L1W and Wytham Woods, while SegmentAnyTree performs slightly better with a coverage of 77.5% and 41.2%. Increasing the amount of training data continuously increases the performance of TreeLearn. Compared to TreeLearn small, the coverage increases by 3% on L1W and 3.1% on Wytham Woods when training with all the available data. In Appendix D, we report the offset losses for all TreeLearn models.
3.3 Detailed instance segmentation analysis
Figure 4: Comparison of instance segmentation results for L1W and Wytham Woods plots. TreeLearn results obtained from large (L1W) or mid setting (Wytham Woods).
@@ 3.3 Detailed instance segmentation analysis
Figure 5: Instance segmentation performance. The top row depicts metrics for segments along the x- and y- axis, while the bottom row depicts metrics for segments along the z-axis. (a) shows results for L1W and (b) for Wytham Woods.
In this section, we take a more detailed look at the quality of the segmentations. In addition to an overall better coverage, TreeLearn achieves a better precision and recall than the comparison methods when it comes to assigning points to trees (Table 3). Irrespective of the method, recall is higher than precision. On Wytham Woods, this is caused mainly by merged trees: If a prediction is a merged tree that consists of two ground truth trees, we match both of the corresponding ground truth trees to this prediction (see Section2.5.2) and recall will be close to 100%. Precision will only be around 50% in this case, as the points of the other tree are also included in the prediction. Wytham Woods has many omission errors due to merged trees (Figure 4a). However, we also observe this pattern on L1W where merged trees are uncommon. We conjecture that, on this dataset, it is caused by the branches of larger trees protruding into smaller trees. If these branches are incorrectly assigned to the smaller tree, the precision of the smaller tree is affected considerably. On the other hand, a falsely assigned branch does not affect the recall of the larger tree to the same degree. The only exception to this pattern of results is SegmentAnyTree on L1W. This is most likely a consequence of the extraordinarily high commission error rate (49.1%) that negatively affects recall.
@@
A more fine-grained understanding can be achieved by evaluating the segmentation performance at different locations of a tree. For this, we partitioned the points of a ground truth tree and the corresponding prediction into subsets based on horizontal distance to the tree trunk and vertical distance to the forest ground as described in Section 2.5.2. Figure 5 shows precision, recall and coverage for the two partitions. For all methods, performance decreases for points farther away from the trunk due to increased interactions with other trees (Figure 5, first row). However, this trend is less pronounced for TreeLearn, indicating a certain stability with respect to the distance to the trunk. This is especially true for models fine-tuned with manually labeled data. Interestingly, the recall of ForAINet and SegmentAnyTree is already far from 100% on both L1W and Wytham Woods for parts that are very close to the trunk, where segmentation is easiest. This can be explained by coarse segmentation mistakes, such as commission errors that make up large parts of the tree crown (e.g. Figure 4c).
For the vertical partition on L1W (Figure 5a, second row), the methods tend to exhibit a visible performance low for intermediate heights, where trees tend to have their maximum width and therefore many interactions with other trees. Furthermore, the area directly above the ground poses difficulties for all methods. Potential explanations for this are low-hanging branches that are hard to clearly assign to a specific tree (also discussed in Section 3.4) or ambiguities between trees and understory. Wytham Woods is structurally different from L1W. It contains many smaller trees whose trunks are often very close to each other (Figure 4a). In higher parts of the forest, interactions with other trees tend to decrease since these smaller trees are not present anymore. This is reflected by the results of the vertical partition (Figure 5b, second row), which indicate an increasing performance from lower to higher.
3.4 Error analysis
Figure 6: Failure cases in tree clustering due to systematic errors in the space of offset-shifted coordinates.
In this section, we examine failure cases of our method by taking a look at the space of offset-shifted coordinates that are used to obtain tree clusters (see Section 2.2.2). Firstly, parts of a tree that cannot be clearly assigned to the ground truth tree base by the model can have a biased offset prediction. If enough points are projected to an incorrect location, the initial clustering algorithm (see Section 2.2.4) might erroneously identify these points as a tree instance. The remaining points are then assigned to this incorrect instance during postprocessing (see Section 2.2.5). This failure case is depicted in Figure 6a.
Another common error arises when the model is unsure as to which tree a point belongs to. For example, this can occur for points located at the boundary of two intersecting tree crowns. In this case, these points are typically projected in-between the two tree bases resulting in a characteristic pattern of sparse connections between dense tree base clusters. If a third tree is close, uncertain points might be erroneously assigned to this tree based on spatial proximity in the space of offset-shifted coordinates. Figure 6b depicts such a failure case. Failures of this kind can lead to oddly looking mistakes where points that are far away from the ground truth tree are falsely assigned to it.
Finally, we will give an explanation for the increased number of commission errors on L1W for the fine-tuned versions of TreeLearn compared to the base version trained on noisy labels. TreeLearn noisy was trained with a dataset where trees smaller than 10 m were labeled as non-tree. In contrast, there are many smaller trees in the manually labeled datasets that have been used for fine-tuning. A qualitative inspection of the commission errors suggests that low-hanging branches might have been erroneously identified as small trees by the fine-tuned models. Points belonging to such branches were not projected to the tree base and consequently formed separate clusters (Figure 6c). This suggests that small trees are a challenge since they introduce edge cases that require additional training data for accurate prediction.
However, simply adding more training data will not address the core challenge of this tree segmentation paradigm: Forests are complex structures, where the identification and separation of individual tree instances is a highly non-trivial task. As a consequence, neural networks trained for offset prediction will often interpolate between two trees or project points to entirely incorrect positions, complicating accurate instance prediction. To improve tree separation, one could incorporate additional parameters, such as tree height, into offset prediction. Furthermore, including some form of regularization into the offset prediction training objective could help produce more clearly separable clusters. However, solving the issues associated with this paradigm will, in our opinion, also require more general computer vision research on how to effectively discern objects with complex instance boundaries, of which trees are a prime example.